OpenStudy (anonymous):

What is the simplified form of (15x^6)/(20y^5) times (6y^2)/(5x^4)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\Huge \frac{15x^6}{20y^5}\cdot\frac{6y^2}{5x^4}$$

OpenStudy (anonymous):

The answer choices are 9x^2/10y^3 9x^3/10y^2 10x^3/9y^2 10x^2/9y^3

OpenStudy (muzzack):

Step by step solution : Step 1 : \(\huge\ (3•5^x6)\) Simplify ———————— \(\huge\ (22•5y^5)\) Canceling out \(\huge\ 5\) as it appears on both sides of the fraction line Equation at the end of step 1 : \(3x^6\) \(6y^2\) ——— • ——— \(4y^5\) \(5x^4\)

OpenStudy (muzzack):

Step 2 : \(3x6\) \(6y2\) Multiply ——— by ——— \(4y5\) \(5x4\) \(x^6\) divided by \(\huge\ x^4\) = \(\huge\ x^{(6 - 4)}\) = \(x^2\) \(y^2~divided~by~y^5 = y^{(2 - 5)} = y^{(-3)} = 1/y^3\) Final result : \(\large\ 9x^2\) ———— \(\large\ 10y^3\)

OpenStudy (anonymous):

Ok, thanks